# MT2ST: Adaptive Multi-Task to Single-Task Learning

### v1

@@ Abstract
In this paper, we propose the Multi-Task to Single-Task (MT2ST) framework, a novel approach that significantly enhances the efficiency and accuracy of word embedding training while concurrently addressing prevalent issues such as overfitting. Our empirical studies demonstrate that MT2ST can reduce training time by 67% when contrasted with single-task learning approaches, and by 13% compared to traditional multi-task learning methods. These findings underscore MT2ST’s potential to be a powerful tools for word embedding training acceleration.
@@ I Introduction
Traditional word embedding training methods predominantly rely on either MTL, which offers a wide learning scope at the expense of task-specific detail, or STL, which excels in focus on specific task but lacks the comprehensiveness of MTL. Our research introduces a novel paradigm, the MT2ST framework, which combines comprehensive scope of MTL and task-sepecific precision of STL together. [CPKC09]
@@ II-A Challenges in Single-Task Word Embedding Training
Computational Complexity: Training involves complex mathematical computations. These algorithms, while effective, require high computational resources, leading to increased training duration and processing power needs.
@@ III-C Model Overview
Figure 1: MT2ST Word Embedding Training Framework Overview [^2] This is a footnote explaining the MT2ST model.
22footnotetext: Code: https://github.com/NoakLiu/MT2ST-Word-Embeddings-Acceleration.
@@ V Discussion
Our empirical experiments reveal that the MT2ST model notably outperforms conventional embedding training methods in terms of training efficiency and overall quality. The newly developed embeddings exhibit superior performance and robustness across a range of downstream NLP tasks, including sentiment analysis, part-of-speech tagging, and named entity recognition. These results underscore the effectiveness of MT2ST in reducing training times and lessening computational resource requirements. This efficiency makes MT2ST a valuable solution for fast training and high-quality word representations.

### v2

@@ Abstract
In this paper, we propose the Multi-Task to Single-Task (MT2ST) framework, a novel approach that significantly enhances the efficiency and accuracy of word embedding training while concurrently addressing prevalent issues such as overfitting. Our empirical studies demonstrate that MT2ST can reduce training time by 67% when contrasted with single-task learning approaches, and by 13% compared to traditional multi-task learning methods. These findings underscore MT2ST’s potential to be a powerful tools for word embedding training acceleration. The code implementation is can be found at: https://github.com/NoakLiu/MT2ST-Word-Embeddings-Acceleration.
@@ I Introduction
Traditional word embedding training methods predominantly rely on either MTL, which offers a wide learning scope at the expense of task-specific detail, or STL, which excels in focus on specific task but lacks the comprehensiveness of MTL. [ZWX+24]. Our research introduces a novel paradigm, the MT2ST framework, which combines comprehensive scope of MTL and task-sepecific precision of STL together. [CPKC09]
@@ II-A Challenges in Single-Task Word Embedding Training
Computational Complexity: Training involves complex mathematical computations. These algorithms, while effective, require high computational resources, leading to increased training duration and processing power needs. Some recent work such as [LWJV24], [LJP24] has proven the importance of efficient computing in reducing computational resources through algorithm and system co-design optimization.
@@ III-C Model Overview
Figure 1: MT2ST Word Embedding Training Framework Overview22footnotemark: 2
@@ V Discussion
Our empirical experiments reveal that the MT2ST model notably outperforms conventional embedding training methods in terms of training efficiency and overall quality. The newly developed embeddings exhibit superior performance and robustness across a range of downstream NLP tasks, including sentiment analysis, part-of-speech tagging, and named entity recognition. These results underscore the effectiveness of MT2ST in reducing training times and lessening computational resource requirements. The multi-task to single task learning also has applications in some recent work, such as [WCL24] [LWC+24] [XQW+24] [DLL24]. This efficiency makes MT2ST a valuable solution for fast training and high-quality word representations.